OpenStudy (anonymous):

Factor Completely: 7h4-4h3+28h2-16h the 4,2, and 2 are superscripts.

DivineSolar (divinesolar):

here you go.

OpenStudy (zehanz):

Just giving away the answer isn't that helpful imo. The first step is clear: there is an h in every term, so: \(7h^4-4h^3+28h^2-16h=h(7h^3-4h^2+28h-16)\). Now we seem to be stuck. How to factor this one? There are four terms! Now wait a minute...why not change the order of the terms: \(h(7h^3-28h-4h^2-16)\). Why? Just because the 7 and 28 seem to be related to each other in the same way the 4 and 16 are: 7*4=28 and 4*4=16.

OpenStudy (zehanz):

(obviously, the second term in the brackets is is +28h). Next step is to factor the four terms in two groups: \(7h^3+28h=7h(h^2+4)\) \(-4h^2-16=-4(h^2+4)\) Can you guess where this is going to?

OpenStudy (zehanz):

@kyliexanne_ : do you understand so far?

OpenStudy (zehanz):

The first factor was \(h\). So now we've found \(h^2+4\) is also a factor. If we put everything together, we get: \(7h^4-4h^3+28h^2-16h=h(h^2+4)(7h-4)\). Done!

OpenStudy (zehanz):

The first factor was \(h\). So now we've found \(h^2+4\) is also a factor. If we put everything together, we get: \(7h^4-4h^3+28h^2-16h=h(h^2+4)(7h-4)\). Done!